A domestic radiator is 2500mm long and 750mm high and has a surface temperature of 86ºC. If the mean room temperature is 18ºC, determine the rate of heat output from the radiator.
Answer #1

Length of the radiator = 2500mm = 2.5m

Height of the radiator = 750mm = 0.7m

Consider the unit width of the radiator , b = 1m

Surface temperature (Th) = 86 degree C

Room Temperature (Tc) = 18 degree C

Consider the material of construction of radiator be aluminium, Thermal conductivityof aluminium = 237W/mK

Q = (Th- Tc) / (b/KA) = (86-18) / (1/237* 1.75) = 28.20KW